A certain elastic conducting material is stretched into a circular loop of 10.0 cm radius. It is placed with its plane perpendicular to a uniform 0.800T magnetic field. When released, the radius of the loop starts to shrink at an instantaneous rate of 75.0cm/s. What emf is induced in the loop at that instant 

Grade:12

1 Answers

Khimraj
6 years ago
Magnitude of Induces EMF will be d∅/dt = d(BA)/dt = d(Bπr²)/dt = 2πBr(dr/dt) = 2π(0.8)(0.10)(0.75) = 0.3768 V.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer